# Robust image classification with multi-modal large language models

### v1

@@ I Introduction
Deep Neural Networks (DNNs) are employed in numerous image recognition tasks [1], such as facial recognition, medical imaging, pedestrian segmentation for autonomous driving, and surveillance. For instance, in medical imaging, DNNs assist in detecting diseases like cancer from X-rays and MRIs [2]. In autonomous driving [3], DNNs help recognize pedestrians and road signs, improving safety and navigation. Their remarkable performance has, therefore, led to widespread integration into various applications, impacting user safety, daily routines, services, and product quality. However, these models are vulnerable to adversarial examples [4, 5, 6]—subtle, often imperceptible perturbations in the input data that can mislead the model into making incorrect predictions or classifications. Adversarial examples are typically generated through gradient-based evasion attacks [7, 8], which exploit the model’s sensitivity to minor input changes to induce erroneous predictions, often with high confidence. Furthermore, adversarial examples pose a significant threat to the reliability of DNNs, especially in user-facing and safety-critical applications. These attacks not only degrade model performance but also introduce serious concerns about compliance with emerging regulatory standards. For instance, frameworks like the AI Act [9] have imposed stricter demands on the reliability and accountability of AI systems. Therefore, as DNNs deployment expands and regulatory requirements tighten, developing robust defenses against adversarial examples is increasingly important for safe and responsible use.
@@
Consequently, various defensive strategies have been proposed in recent years to address these challenges. Prominent among them are adversarial training and detectors, complementary approaches both representing proactive defenses that align with the security-by-design principle by incorporating security measures during model development [10]. Adversarial training [11, 12] strengthens models by injecting knowledge of adversarial attacks during their training (e.g., including adversarial examples in the training data), improving their resilience to future attacks. Detector-based defenses, on the other hand, encompass techniques that detect and mitigate/reject attacks as they occur. Examples include anomaly detection systems that reject potentially dangerous inputs in real-time and abstain from predicting them [13, 14, 15]. However, each of these defense strategies has its own limitations. Adversarial training, while effective, can be complex to develop and resource-intensive [16], requiring continuous adaptation or retraining to address evolving attacks. Conversely, detectors rely on recognizing deviations from normal behavior, which may not be effective against new attacks [10].
@@
To address these challenges, we propose Multi-Shield, a solution that integrates the strengths of both adversarial training and detection defenses while handling multi-modal information. The idea is to combine adversarial training defenses with a rejection module that incorporates both visual and textual information during prediction. In particular, Multi-Shield uses an adversarially trained model and a multi-modal model (i.e., CLIP [17]), which analyzes visual and textual patterns to verify their alignment. Specifically, the multi-modal model uses visual images and textual descriptions related to the classes in the dataset to assess semantic agreement, employing zero-shot predictions. As illustrated in Figure 1, when there is an agreement between the proactively trained image classifier and the multi-modal model, the system provides a prediction. However, if there is a discrepancy, the system abstains from responding, as the lack of agreement suggests potential adversarial manipulation or uncertainty. Our experiments demonstrate the effectiveness of Multi-Shield on two benchmark datasets, CIFAR-10 and ImageNet. We applied Multi-Shield to both non-robust models and robust models ($6$ for CIFAR-10 and $4$ for ImageNet) developed with adversarial training. The results reveal that Multi-Shield significantly increase robustness for proactively trained models, as it successfully rejects adversarial examples that would otherwise bypass them. Overall, these results highlight Multi-Shield ability to complement existing defenses and provide an additional layer of security for them.
@@
  • we propose Multi-Shield, a novel defense mechanism that leverages multi-modal data, specifically visual and textual inputs, to identify and reject adversarial examples;
  • we explore the interaction between adversarial training and a multi-modal detector, demonstrating how Multi-Shield can be effectively integrated as a supplementary layer to enhance overall model robustness;
  • we assess the robustness improvements provided by multi-modal information against adversarial attacks by evaluating Multi-Shield on two vision datasets and ten models.
The rest of the paper is structured as follows: Section II provides an overview of adversarial defenses and the principles of multi-modal large language models. In Section III, we present the architecture and inner workings of Multi-Shield. Section IV outlines our experimental setup and results, and Section V concludes with limitations and future research directions.
@@ II Background and Related Work
In this section, we present background information and review related work on adversarial attacks and defenses, as well as multi-modal large language models.
@@ II-A Adversarial examples
where $L(\cdot)$ in Eq. (missing) 1 is a loss function that penalizes correct classification of $\bm{x}^{\prime}$ by $f$. A commonly used loss function is the Difference of Logits [6], expressed as:
@@
with $f_{i}(\bm{x})$ denoting the logit score of $\bm{x}$ associated with class $i$. Under this loss, the optimization program aims to find an adversarial example $\bm{x}^{\prime}$ that decreases the logit corresponding to the true class label $y$ while increasing the logits of competing incorrect classes. As a result, the model is driven to misclassify the input by predicting the incorrect class with the highest logit score. The constraint in Eq. (missing) 2 ensures that the adversarial example $\bm{x}^{\prime}$ remains similar to the original input $\bm{x}$ by limiting the norm of their difference, with typical choices for $p\in\{0,1,2,\infty\}$ [6]. Lastly, the constraint in Eq. (missing) 3 ensures that $\bm{x}^{\prime}$ remains within the valid input space $[0,1]^{d}$, a typical requirement for image recognition tasks. Given the above formulation, several optimization methods have been proposed to solve this problem, including PGD [18], FGSM [11], and many others [8]. Among them, AutoAttack [7] stands out as a state-of-the-art approach due to its ability to generate adversarial examples effectively without parameter tuning, making it well-suited for reliable and consistent benchmarking of model robustness.
@@ II-B Adversarial Defenses
Prominent proactive defenses against adversarial examples in DNNs fall into two main categories: robust optimization techniques and adversarial detectors [10, 15]. Among the most effective methods in the first category is adversarial training [11, 12], which strengthens robustness by exposing the model to adversarially perturbed inputs during the training process and has consistently proven to be one of the most effective defense mechanisms to date [12, 7]. Detectors, on the other hand, focus on mechanisms that enable the model to abstain from making predictions when it suspects inputs to be adversarial. The pioneering approach by [19] involves rejecting samples whose feature representations are significantly distant from class centroids. Similarly, [14] propose a distance-based rejection method using thresholds applied to outputs of a multi-class SVM to estimate data feature representation distribution. Later works extend the inspection of feature representations beyond the final layer of the network. [13] leverage late-stage ReLU activations combined with a quantized RBF-SVM to detect adversarial examples, though this method requires adversarial training, increasing computational cost. Alternatively, [20] present a KNN-based method that evaluates intermediate layer representations and rejects inputs with inconsistencies, though it incurs substantial overhead due to the need for distance computations with respect to training data at each layer. [15] propose a method for evaluating the consistency of feature representations across different layers, using a threshold for distance-based rejection.
@@
II-C Multi-modal Vision-Language Models
Multi-modal models represent a recent development in AI, enabling the simultaneous processing of multiple types of data [21]. These models are trained to process data from different modalities, identifying patterns aligning the information. In particular, vision-language map images and textual descriptions into a shared semantic space, facilitating tasks that require a joint understanding of both modalities [22]. Applications of these models include image captioning [23], visual question answering [24], and text-to-image retrieval [25], where the combination of visual and textual information allows models to better capture contextual meaning and improve performance. In this work, we focus on vision-language models applied to image recognition tasks. Several prominent vision-text multi-modal models have emerged, including VisualBERT [26], VilBERT [27], and VisionLLaMA [28], each designed to capture complex relationships between visual and textual inputs. These models typically employ transformer-based architectures [29] to encode visual and textual data into a unified representation. Another notable example of vision-text multi-modal is CLIP (Contrastive Language-Image Pretraining) [17], which aligns visual and textual embeddings using large-scale datasets of paired images and captions. CLIP works by training on a contrastive learning objective, ensuring that semantically similar images and text are positioned closely within a shared embedding space, leading to strong performance in tasks such as zero-shot image classification and text-to-image retrieval [17].
@@ III-A Multishield’s Workflow
The Multi-Shield framework operates in three distinct phases, depicted in Figure 1, each contributing to a robust image classification process that leverages both unimodal and multi-modal analysis. We name these phases with: unimodal prediction, multi-modal alignment, and Multi-Shield decision
@@
Multi-modal Alignment. In this second phase, Multi-Shield uses the CLIP model as a zero-shot vision-language classifier [17]. It compares the visual representation of the input image with class description prompts to determine alignment. For each class, we first create natural language prompts in the format ‘Picture of a [object]’ (Step 3). For example, for classes such as ‘parrot’ and ‘dog,’ the prompts would be ‘Picture of a parrot’ and ‘Picture of a dog’, respectively. The text prompts serve as a simplified but effective method to represent each class conceptually; however, they could be further enriched by including additional descriptive characteristics (e.g., mentioning that parrots have feathers and wings). During prediction, these prompts, along with the input image, are processed through CLIP’s dual encoders: one for images and one for text (Step 4). This process generates visual and textual embeddings that capture the features of both modalities, with CLIP aiming to maximize the alignment between the image and the textual description that best matches it. Finally, Multi-Shield computes the alignment scores using cosine similarity, selecting the class with the highest score as the final prediction from the CLIP model (Step 5).
Multi-Shield Decision. In the final phase, Multi-Shield compares the predictions from the unimodal image classifier and the multi-modal CLIP model (Step 6), each capturing distinct patterns in the data. If both models agree, Multi-Shield outputs their shared prediction; if not, it abstains (Step 7). As shown on the left side of Figure 1, for a clean image, both models confidently identify the class ‘parrot,’ leading Multi-Shield to output the agreed-upon prediction. Conversely, on the right side of Figure 1, adversarial noise causes the unimodal classifier to misclassify the image, while the CLIP model still correctly aligns the image with the ‘parrot’ class. This disagreement drives Multi-Shield to abstain, acknowledging that the adversarial input may have misled one model. Unlike typical robust models, which do not employ a rejection mechanism, Multi-Shield introduces this feature to abstain when uncertain (e.g., input is suspected to be adversarial) and ensures predictions are made only when both models agree.
@@ III-B Problem Formulation
Here, $R(\bm{x})$ represents the level of agreement between the unimodal image classifier and the zero-shot multi-modal prediction from CLIP, serving as a score to gauge the model’s uncertainty about the input. When $R(\bm{x})$ is non-positive, it indicates that the image classifier and CLIP are in agreement, suggesting high confidence in the prediction. Conversely, if $R(\bm{x})$ is positive, the model abstains from making a prediction due to uncertainty. The rejection score $R(\bm{x})$ of Multi-Shield is computed as:
@@
where $h_{\rm{img}}(\bm{x})\in\mathbb{R}^{d}$ is the visual embedding of the input image and $h_{\rm{txt}}(P_{i})\in\mathbb{R}^{d}$ is the textual embedding for the class prompt $i$, both generated using the encoders of the multi-modal model. Multi-Shield evaluates the alignment score for each class prompt and picks the highest one to perform zero-shot classification using the CLIP model. Finally, when Multi-Shield abstains, the positive score in $R(\bm{x})$ indicates the level of disagreement between the image classifier and CLIP’s classification, supporting decisions on abstention reliability and subsequent input handling.
@@ III-C Attacking Multi-Shield
To evaluate the worst-case robustness of Multi-Shield, we assess its performance against an adaptive attacker [30], who is fully aware of the defense and actively seeks to circumvent it. Adaptive attacks indeed provide a more rigorous and stronger test of defenses [20, 31, 15]. The attacker, having access to the image recognition model and the rejection mechanism in Multi-Shield, aims to simultaneously deceive the unimodal image classifier and manipulate the zero-shot multi-modal classifier, causing both to agree on an incorrect label. The adaptive attack can thus be formalized by revisiting the optimization program in Eq. (missing) 1-Eq. (missing) 3 as follows:
@@ IV Experiments
TABLE I: Experimental results on baseline and adversarially pre-trained models (6 on CIFAR-10, 4 on ImageNet) evaluated using clean accuracy (%), robust accuracy (%), and rejection ratio (%). The defense is assessed using AutoAttack with perturbation size $\varepsilon=8/255$ across three defense scenarios: without Multi-Shield (Baseline), with Multi-Shield (Multi-Shield), and adaptively attacking Multi-Shield (Multi-Shield - Adaptive Attack). Note that performing a standalone attack on either of the underlying clip models yields a robust accuracy of 0%.

### v2

@@ I Introduction
Deep Neural Networks (DNNs) are employed in numerous image recognition tasks [Rawat and Wang, 2017], such as facial recognition, medical imaging, pedestrian segmentation for autonomous driving, and surveillance. For instance, in medical imaging, DNNs assist in detecting diseases like cancer from X-rays and MRIs [Mall et al., 2023]. In autonomous driving [Yurtsever et al., 2020], DNNs help recognize pedestrians and road signs, improving safety and navigation. Their remarkable performance has, therefore, led to widespread integration into various applications, impacting user safety, daily routines, services, and product quality. However, these models are vulnerable to adversarial examples [Biggio et al., 2013, Szegedy et al., 2014, Carlini and Wagner, 2017a]—subtle, often imperceptible perturbations in the input data that can mislead the model into making incorrect predictions or classifications. Adversarial examples are typically generated through gradient-based evasion attacks Croce and Hein [2020], Cinà et al. [2025a, b], which exploit the model’s sensitivity to minor input changes to induce erroneous predictions, often with high confidence. Furthermore, adversarial examples pose a significant threat to the reliability of DNNs, especially in user-facing and safety-critical applications. These attacks not only degrade model performance but also introduce serious concerns about compliance with emerging regulatory standards. For instance, frameworks like the AI Act [Madiega, 2021] have imposed stricter demands on the reliability and accountability of AI systems. Therefore, as DNNs deployment expands and regulatory requirements tighten, developing robust defenses against adversarial examples is increasingly important for safe and responsible use.
@@
Consequently, various defensive strategies have been proposed in recent years to address these challenges. Prominent among them are adversarial training and detectors, complementary approaches both representing proactive defenses that align with the security-by-design principle by incorporating security measures during model development [Biggio and Roli, 2018]. Adversarial training [Goodfellow et al., 2015, Madry, 2017] strengthens models by injecting knowledge of adversarial attacks during their training (e.g., including adversarial examples in the training data), improving their resilience to future attacks. Detector-based defenses, on the other hand, encompass techniques that detect and mitigate/reject attacks as they occur. Examples include anomaly detection systems that reject potentially dangerous inputs in real-time and abstain from predicting them [Lu et al., 2017, Melis et al., 2017, Sotgiu et al., 2020]. However, each of these defense strategies has its own limitations. Adversarial training, while effective, can be complex to develop and resource-intensive [Shafahi et al., 2019], requiring continuous adaptation or retraining to address evolving attacks. Conversely, detectors rely on recognizing deviations from normal behavior, which may not be effective against new attacks [Biggio and Roli, 2018].
@@
To address these challenges, we propose Multi-Shield, a solution that integrates the strengths of both adversarial training and detection defenses while handling multimodal information. The idea is to combine adversarial training defenses with a rejection module that incorporates both visual and textual information during prediction. In particular, Multi-Shield uses an adversarially trained model and a multimodal model (i.e., CLIP [Radford et al., 2021]), which analyzes visual and textual patterns to verify their alignment. Specifically, the multimodal model uses visual images and textual descriptions related to the classes in the dataset to assess semantic agreement, employing zero-shot predictions. As illustrated in Figure 1, when there is an agreement between the proactively trained image classifier and the multimodal model, the system provides a prediction. However, if there is a discrepancy, the system abstains from responding, as the lack of agreement suggests potential adversarial manipulation or uncertainty. Our experiments demonstrate the effectiveness of Multi-Shield on six benchmark datasets (i.e., CIFAR-10, ImageNet, Caltech-101, Food-101, Oxford-IIIT pets, STL-10). We applied Multi-Shield to both non-robust models and robust models developed with adversarial training. The results reveal that Multi-Shield significantly increases robustness for proactively trained models, as it successfully rejects adversarial examples that would otherwise bypass them. Overall, these results highlight Multi-Shield’s ability to complement existing defenses and provide an additional layer of security for them.
@@
  • we propose Multi-Shield, a novel defense mechanism that leverages multimodal data, specifically visual and textual inputs, to identify and reject adversarial examples;
  • we explore the interaction between adversarial training and a multimodal detector, demonstrating how Multi-Shield can be effectively integrated as a supplementary layer to enhance overall model robustness;
  • we assess the robustness improvements provided by multimodal information against adversarial attacks by evaluating Multi-Shield on six datasets and thirteen models.
The rest of the paper is structured as follows: Section II provides an overview of adversarial defenses and the principles of multimodal large language models. In Section III, we present the architecture and inner workings of Multi-Shield. Section IV outlines our experimental setup and results, and Section V concludes with limitations and future research directions.
@@ II Background and Related Work
In this section, we present background information and review related work on adversarial attacks and defenses, as well as multimodal large language models.
@@ II-A Adversarial examples
where $L(\cdot)$ in Eq. (missing) 1 is a loss function that penalizes correct classification of $\bm{x}^{\prime}$ by $f$. A commonly used loss function is the Difference of Logits [Carlini and Wagner, 2017a], expressed as:
@@
with $f_{i}(\bm{x})$ denoting the logit score of $\bm{x}$ associated with class $i$. Under this loss, the optimization program aims to find an adversarial example $\bm{x}^{\prime}$ that decreases the logit corresponding to the true class label $y$ while increasing the logits of competing incorrect classes. As a result, the model is driven to misclassify the input by predicting the incorrect class with the highest logit score. The constraint in Eq. (missing) 2 ensures that the adversarial example $\bm{x}^{\prime}$ remains similar to the original input $\bm{x}$ by limiting the norm of their difference, with typical choices for $p\in\{0,1,2,\infty\}$ [Carlini and Wagner, 2017a]. Lastly, the constraint in Eq. (missing) 3 ensures that $\bm{x}^{\prime}$ remains within the valid input space $[0,1]^{d}$, a typical requirement for image recognition tasks. Given the above formulation, several optimization methods have been proposed to solve this problem, including PGD [Madry et al., 2018], FGSM [Goodfellow et al., 2015], and many others [Cinà et al., 2025b]. Among them, AutoAttack [Croce and Hein, 2020] stands out as a state-of-the-art approach due to its ability to generate adversarial examples effectively without parameter tuning, making it well-suited for reliable and consistent benchmarking of model robustness.
@@ II-B Adversarial Defenses
Prominent proactive defenses against adversarial examples in DNNs fall into two main categories: robust optimization techniques and adversarial detectors [Biggio and Roli, 2018, Sotgiu et al., 2020]. Among the most effective methods in the first category is adversarial training [Goodfellow et al., 2015, Madry, 2017, Li et al., 2024], which strengthens robustness by exposing the model to adversarially perturbed inputs during the training process and has consistently proven to be one of the most effective defense mechanisms to date [Madry, 2017, Croce and Hein, 2020]. Detectors, on the other hand, focus on mechanisms that enable the model to abstain from making predictions when it suspects inputs to be adversarial. The pioneering approach by Bendale and Boult [2016] involves rejecting samples whose feature representations are significantly distant from class centroids. Similarly, Melis et al. [2017] propose a distance-based rejection method using thresholds applied to outputs of a multi-class SVM to estimate data feature representation distribution. Later works extend the inspection of feature representations beyond the final layer of the network. Lu et al. [2017] leverage late-stage ReLU activations combined with a quantized RBF-SVM to detect adversarial examples, though this method requires adversarial training, increasing computational cost. Alternatively, Papernot and McDaniel [2018] present a KNN-based method that evaluates intermediate layer representations and rejects inputs with inconsistencies, though it incurs substantial overhead due to the need for distance computations with respect to training data at each layer. Sotgiu et al. [2020] propose a method for evaluating the consistency of feature representations across different layers, using a threshold for distance-based rejection.
@@
II-C Multimodal Vision-Language Models
Multimodal models represent a recent development in AI, enabling the simultaneous processing of multiple types of data [Yin et al., 2023]. These models are trained to process data from different modalities, identifying patterns aligning the information. In particular, vision-language map images and textual descriptions into a shared semantic space, facilitating tasks that require a joint understanding of both modalities [Du et al., 2022]. Applications of these models include image captioning [Hossain et al., 2019], visual question answering [Wu et al., 2017], and text-to-image retrieval [Cao et al., 2022], where the combination of visual and textual information allows models to better capture contextual meaning and improve performance. In this work, we focus on vision-language models applied to image recognition tasks. Several prominent vision-text multimodal models have emerged, including VisualBERT [Li et al., 2019], VilBERT [Lu et al., 2019], and VisionLLaMA [Chu et al., 2024], each designed to capture complex relationships between visual and textual inputs. These models typically employ transformer-based architectures [Vaswani et al., 2017] to encode visual and textual data into a unified representation. Another notable example of vision-text multimodal is CLIP (Contrastive Language-Image Pretraining) [Radford et al., 2021], which aligns visual and textual embeddings using large-scale datasets of paired images and captions. CLIP works by training on a contrastive learning objective, ensuring that semantically similar images and text are positioned closely within a shared embedding space, leading to strong performance in tasks such as zero-shot image classification and text-to-image retrieval [Radford et al., 2021].
@@ III-A Multishield’s Workflow
The Multi-Shield framework operates in three distinct phases, depicted in Figure 1, each contributing to a robust image classification process that leverages both unimodal and multimodal analysis. We name these phases with: unimodal prediction, multimodal alignment, and Multi-Shield decision
@@
Multimodal Alignment. In this second phase, Multi-Shield uses the CLIP model as a zero-shot vision-language classifier [Radford et al., 2021]. It compares the visual representation of the input image with class description prompts to determine alignment. For each class, we automatically create natural language prompts in the format ‘Picture of a [object]’ (Step 3). For example, for classes such as ‘parrot’ and ‘dog,’ the prompts would be ‘Picture of a parrot’ and ‘Picture of a dog’, respectively. The text prompts serve as a simplified but effective method to represent each class conceptually; however, with some human effort, they could be further enriched by including additional descriptive characteristics (e.g., mentioning that parrots have feathers and wings). During prediction, these prompts, along with the input image, are processed through CLIP’s dual encoders: one for images and one for text (Step 4). This process generates visual and textual embeddings that capture the features of both modalities, with CLIP aiming to maximize the alignment between the image and the textual description that best matches it. Finally, Multi-Shield performs zero-shot classification with CLIP by computing alignment scores using cosine similarity and selecting the class with the highest score as the final prediction [Radford et al., 2021] (Step 5).
Multi-Shield Decision. In the final phase, Multi-Shield compares the predictions from the unimodal image classifier and the multimodal CLIP model (Step 6), each capturing distinct patterns in the data. If both models agree, Multi-Shield outputs their shared prediction; if not, it abstains (Step 7). As shown on the left side of Figure 1, for a clean image, both models confidently identify the class ‘parrot,’ leading Multi-Shield to output the agreed-upon prediction. Conversely, on the right side of Figure 1, adversarial noise causes the unimodal classifier to misclassify the image, while the CLIP model still correctly aligns the image with the ‘parrot’ class. This disagreement drives Multi-Shield to abstain, acknowledging that the adversarial input may have misled one model. Unlike typical robust models, which do not employ a rejection mechanism, Multi-Shield introduces this feature to abstain when uncertain (e.g., input is suspected to be adversarial) and ensures predictions are made only when both models agree.
@@ III-B Problem Formulation
Here, $R(\bm{x})$ represents the level of agreement between the unimodal image classifier and the zero-shot multimodal prediction from CLIP, serving as a score to gauge the model’s uncertainty about the input. When $R(\bm{x})$ is non-positive, it indicates that the image classifier and CLIP are in agreement, suggesting high confidence in the prediction. Conversely, if $R(\bm{x})$ is positive, the model abstains from making a prediction due to uncertainty. The rejection score $R(\bm{x})$ of Multi-Shield is computed as:
@@
where $h_{\rm{img}}(\bm{x})\in\mathbb{R}^{d}$ is the visual embedding of the input image and $h_{\rm{txt}}(P_{i})\in\mathbb{R}^{d}$ is the textual embedding for the class prompt $i$, both generated using the encoders of the multimodal model. Multi-Shield evaluates the alignment score for each class prompt and picks the highest one to perform zero-shot classification using the CLIP model. Finally, when Multi-Shield abstains, the positive score in $R(\bm{x})$ indicates the level of disagreement between the image classifier and CLIP’s classification, supporting decisions on abstention reliability and subsequent input handling.
@@ III-C Attacking Multi-Shield
To evaluate the worst-case robustness of Multi-Shield, we assess its performance against an adaptive attacker [Carlini and Wagner, 2017b], who is fully aware of the defense and actively seeks to circumvent it. Adaptive attacks indeed provide a more rigorous and stronger test of defenses [Papernot and McDaniel, 2018, Carlini et al., 2019, Sotgiu et al., 2020]. The attacker, having access to the image recognition model and the rejection mechanism in Multi-Shield, aims to simultaneously deceive the unimodal image classifier and manipulate the zero-shot multimodal classifier, causing both to agree on an incorrect label. The adaptive attack can thus be formalized by revisiting the optimization program in Eq. (missing) 1-Eq. (missing) 3 as follows:
@@ IV Experiments
TABLE I: Experimental results on baseline and adversarially pre-trained models (6 on CIFAR-10, 4 on ImageNet, and 3 on Caltech-101, Food-101, Oxford-IIIT Pets, and STL-10) evaluated using clean accuracy (%), robust accuracy (%), and rejection ratio (%). The defense is assessed using AutoAttack with perturbation size $\varepsilon=8/255$ across three defense scenarios: without Multi-Shield (Baseline), with Multi-Shield (Multi-Shield), and adaptively attacking Multi-Shield (Multi-Shield - Adaptive Attack). Note that performing the attack on either of the underlying clip models yields a robust accuracy of 0%.